# Understanding and Mitigating Miscalibration in Prompt Tuning for Vision-Language Models

### v1

@@ Abstract
Confidence calibration is critical for the safe deployment of machine learning models in the real world. However, such issue in vision-language models like CLIP, particularly after fine-tuning, has not been fully addressed. In this work, we demonstrate that existing prompt tuning methods usually lead to a trade-off of calibration between base and new classes: the cross-entropy loss in CoOp causes overconfidence in new classes by increasing textual label divergence, whereas the regularization of KgCoOp maintains the confidence level but results in underconfidence in base classes due to the improved accuracy. Inspired by the observations, we introduce Dynamic Outlier Regularization (DOR) to ensure the confidence calibration on both base and new classes after fine-tuning. In particular, we propose to minimize the feature deviation of novel textual labels (instead of base classes) sampled from a large vocabulary. In effect, DOR prevents the increase in textual divergence for new labels while easing restrictions on base classes. Extensive experiments demonstrate that DOR can enhance the calibration performance of current fine-tuning methods on base and new classes.
@@ 1 Introduction
Large pre-trained vision-language models (VLMs) like CLIP (Radford et al., 2021) have become the de facto standard in today’s zero-shot tasks including image recognition (Wortsman et al., 2022), open-vocabulary segmentation (Liang et al., 2023) and knowledge-augmented retrieval (Ming & Li, 2024). To transfer pre-trained CLIP knowledge to domain-specific downstream tasks efficiently, various parameter-efficient fine-tuning (PEFT) techniques including prompt tuning (Zhou et al., 2022b) and adapter (Gao et al., 2024) have been proposed.  Despite the promising improvement in accuracy, the reliability issue such as confidence calibration in fine-tuned VLMs has been largely overlooked. Without fully understanding the miscalibration in fine-tuned VLMs, it can exacerbate safety concerns in high-stakes applications like medical diagnosis and autonomous driving.
@@
In confidence calibration, we generally expect the model’s confidence level to be consistent with its empirical accuracy. In the literature, zero-shot CLIP is often recognized for its excellent performance in confidence calibration (Minderer et al., 2021). Prior work (Wang et al., 2024) finds that CLIP fined-tuned on base classes generally suffers from miscalibration on novel classes within the same task, where the model is expected to generalize  (Zhou et al., 2022b; Yao et al., 2023). However, they concentrate on novel classes that are not present in the fine-tuning, without adequately explaining the miscalibration issue on base classes. The community still lacks a comprehensive understanding of the fundamental cause and mitigation strategies of the miscalibration issue during fine-tuning.
@@
In this work, we first investigate how current prompt tuning methods (e.g., CoOp (Zhou et al., 2022b), KgCoOp (Yao et al., 2023)) interfere with the calibration of CLIP. We empirically find that existing prompt tuning methods fail to maintain the calibration performance on both base and new classes simultaneously, compromising one of them: CoOp exhibits overconfidence on new classes and KgCoOp provides underconfident predictions on base classes. We provide a thorough explanation from the perspective of textual label divergence. In particular, CoOp increases the divergence of textual label distribution through cross-entropy loss, resulting in excessively high confidence misaligned with actual accuracy on new classes. Instead, KgCoOp hinders the increase of textual distribution divergence, maintaining the confidence level on base and new classes. However, this leads to the underconfidence issue due to the improved accuracy on base classes. There arises a question: Is it possible to ensure the calibration on both base and new classes after fine-tuning?
@@
To tackle the above challenges, we introduce Dynamic Outlier Regularization (DOR). The high-level idea behind DOR is to control the divergence of unseen textual distribution under the supervision of zero-shot CLIP without affecting the vanilla fine-tuning objectives. First, we construct a set of textual outliers from a large lexical database – WordNet (Miller, 1995), and ensure that the selected classes are relevant but non-overlapped with the base classes in the fine-tuning task. In fine-tuning, we reduce the feature discrepancy of novel textual labels between the fine-tuned model and the zero-shot CLIP. In each epoch, the novel textual labels are dynamically sampled from the constructed set of textual outliers. Leveraging dynamic textual outliers, DOR prevents the increase in textual divergence for new labels while easing restrictions on base classes.
@@
We verify the effectiveness of DOR over 11 image classification datasets and four types of ImageNets with covariant shifts, under the evaluation protocol of base-to-new generalization and domain generalization, respectively. Empirical results show that DOR can enhance the overall calibration of existing prompt-tuning methods (see Table 1), on both base and new classes. Moreover, our method can maintain and even improve the generalization performance of those tuning methods (See Table 1 & 2 ). DOR also achieves significant improvements in the presence of covariate shifts. In addition to prompt-based tuning methods, we demonstrate that such a regularization criterion can be extended to visual fine-tuning methods with image outliers.
@@
    We provide an in-depth analysis of textual distribution divergence to understand the miscalibration in fine-tuned CLIP. We also show that current prompt-tuning methods typically lead to a trade-off between base and new classes, compromising one of them.
@@
    We propose DOR, a simple yet effective regularization that ensures the calibration performance on both base and new classes. Our method is compatible with existing prompt-tuning methods and can be extended to visual fine-tuning methods with image outliers.
@@
    We conduct extensive experiments and show that DOR achieves superior performance on a wide range of real-world tasks. For instance, DOR achieves an average of 8.09% reduction in Expected Calibration Error (ECE) for CoOp over the 11 downstream datasets.
@@ 2 Preliminaries
CLIP is a visual-language model that enables to measure the alignment between images and texts (Radford et al., 2021). Recently, CLIP has shown great potential in zero-shot inference for arbitrary classes. Let $\phi:\bm{x}\rightarrow\mathbb{R}^{d}$ and $\psi:\bm{t}\rightarrow\mathbb{R}^{d}$ denote CLIP’s image and text encoders, respectively. Given an image instance $\bm{x}$ and a text label $c$, the logit function of CLIP can be formulated as:
$L_{c}^{clip}\left(\bm{x}_{i}\right)=\tau\cdot\operatorname{sim}\left(\phi(\bm{x}),\psi(\bm{t}_{c})\right).$ (1)
Here $\bm{t}_{c}$ is derived from a hand-crafted prompt like “a photo of a {class}”, where the “{class}” is filled with the text label $c$. $\tau$ is generally set as a pre-trained constant of 100.
@@
$c^{*}=\underset{c\in\mathcal{C}}{\arg\max}\ p(c|\bm{x})=\underset{c\in\mathcal{C}}{\arg\max}\ \frac{e^{L_{c}^{clip}\left(\bm{x}\right)}}{\sum_{i=1}^{C}e^{L_{i}^{clip}\left(\bm{x}\right)}}$ (2)
where $p(c|\bm{x})$ is the predicted probability of class $c$ for the instance $\bm{x}$.
@@
To strengthen the performance of CLIP in downstream applications, prompt tuning methods have been proposed to efficiently fine-tune CLIP on datasets of interest (Zhou et al., 2022a; b; Yao et al., 2023). In particular, prompt tuning optimizes the context prompt only, without retraining the model and updating its weights. For example, CoOp (Zhou et al., 2022b) replaces the hand-crafted textual tokens with a set of learnable textual token $\mathcal{T}=\left\{\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{M}\right\}$, where $M$ is the length of tokens. Thus, the output of fine-tuned CLIP is: $L_{c}^{coop}\left(\bm{x}\right)=\tau\cdot\operatorname{sim}\left(\phi(\bm{x}),\psi(\bm{t}^{\prime}_{c})\right),$ where $\bm{t}_{c}^{\prime}=\left[\bm{v}_{1},\bm{v}_{2},\bm{v}_{3},...\bm{v}_{M},\bm{c}\right]$ and $\bm{c}$ denotes the textual embedding of class $c$. Using a few labeled samples $\mathcal{D}_{\text{ft}}=\{(\bm{x}_{i},c_{i})\}^{N}_{i=1}$, the learnable textual tokens $\mathcal{T}$ are optimized to minimize the cross-entropy (CE) loss $\ell_{ce}$. We refer to the classes used in fine-tuning as base classes, and the remaining labels within the same task as new or novel classes.
In addition to base classes, we generally expect the fine-tuned CLIP to generalize to those new classes within the task. To enhance the generalization ability of the learnable prompt for unseen classes, KgCoOp (Yao et al., 2023) introduces a regularization term to align the learned prompt to the hand-crafted prompt. The optimization of KgCoOp is:
$\mathcal{T}^{\star}=\underset{\mathcal{T}}{\arg\min}\left\{\frac{1}{N}\sum_{i=1}^{N}\ell_{ce}\left(p\left(c_{i}\mid\bm{x}_{i}\right)\right)+\lambda\cdot\frac{1}{C}\sum_{c=1}^{C}\operatorname{sim}\left(\psi\left(\bm{t}_{c}^{\prime}\right),\psi\left(\bm{t}_{c}\right)\right)\right\}.$ (3)
Here, the first term is the cross-entropy loss used in CoOp and the hyperparameter $\lambda$ is used to control the weight of regularization. With $\lambda$ = 0, KgCoOp is degraded to the original CoOp.
@@ 3.1 Empirical study on CLIP calibration
(a) Base
(b) New
Figure 1: Reliability diagram of fine-tuned CLIP (ViT-B/16) on StanfordCars, using prompt tuning methods, CoOp and KgCoOp. ECE: Expected Calibration Error (lower is better). Miscalibration is depicted in pink for overconfidence and purple for underconfidence.
Setup.
To analyze the effects of fine-tuning on CLIP calibration, we first empirically study the calibration performance of fine-tuned VLMs. We use ViT-B-16 pre-trained by OpenAI (Radford et al., 2021) as the zero-shot classification model. In particular, we compare the zero-shot CLIP with two prompt-based tuning methods including CoOp (Zhou et al., 2022b) and KgCoOp (Yao et al., 2023) on StanfordCars dataset (Krause et al., 2013). We evaluate the fine-tuned CLIP under base-to-new protocol: the dataset is split into base and new classes. The model is trained only on a few examples from the base classes and evaluated on examples from both base and new classes.
Prompt tuning leads to a tradeoff between base and new classes
Figure 1 illustrates the calibration performance of zero-shot CLIP, CoOp and KgCoOp on base and new classes. The results show that zero-shot CLIP achieves almost perfect calibration on all classes, while the fine-tuned models cannot maintain the calibration on base and new classes simultaneously, compromising one of them. In particular, CoOp maintains the excellent calibration on base classes but exhibits overconfidence on new classes. Instead, KgCoOp provides underconfident predictions on base classes while preserving the calibration on new classes. This motivates us to further investigate the fundamental cause of miscalibration occurring after fine-tuning.
(a) Confidence
(b) FD score
Figure 2: Results of zero-shot and fined-tuned CLIPs using different prompt tuning methods (KgCoOp and CoOp) on UCF101 dataset.
@@
3.2 Understanding the miscalibration in Fine-tuned CLIP
@@
Definition 1 (Feature Divergence).
@@
where $\mathcal{N}_{M}(\mathbf{z}_{i})$ denotes the set of $M$ nearest neighbors of $\mathbf{z}_{i}$ and $\operatorname{dist}\left(\cdot,\cdot\right)$ is a distance metric like cosine similarity. By averaging these similarity scores across all features, we obtain the overall FD score of a given feature set $\mathrm{FD}(\mathcal{Z})=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, which can represent the divergence of the textual distribution. To investigate the miscalibration caused by CoOp and KgCoOp, we vary the hyperparameter $\lambda$ in Eq. (3). In KgCoOp, $\lambda$ is set to 8.0, and it degenerates to CoOp when $\lambda=0$. We conduct the experiments on UCF101 (Soomro et al., 2012). We present the results in Figure 3.
@@
CoOp leads to overconfidence on new classes by increasing the textual divergence. In the analysis of Subsection 3.1 and Figure 2(a), we show that CLIP tuned by CoOp exhibits overconfidence on new classes, but keeps excellent calibration on base classes. This is caused by the CE loss, which maximizes the posterior $p(y\mid\bm{x})$ for the ground-truth label $y$ and minimizes the probability for other labels. In other words, CE loss tends to enlarge the distance between the image feature and all textual features except the ground truth. As the image feature remains unchanged during fine-tuning, it can be translated to large distances among all textual labels, including base and new classes. This is supported by the results presented in Figure 2(b), which shows that CoOp significantly increases the FD score of textual features compared to the zero-shot CLIP. Consequently, as is shown in Figure 3, the gap between the maximum logit and the others widens for both base and new classes. Therefore, the tuned CLIP with CE loss will make softmax predictions with high confidence on both base and new classes. It aligns with the improved accuracy on base classes, but is not consistent with the nearly unchanged accuracy on new classes. This explains why CLIP tuned by CoOp tends to be overconfident on new classes.
@@
KgCoOp anchors the confidence level by hindering the increase of textual divergence. In the previous analysis, KgCoOp leads to underconfident predictions on base classes while preserving the calibration on new classes. In Figure 2(b), we illustrate the FD scores of textual labels from the zero-shot CLIP and the tuned CLIP by KgCoOp with various $\lambda$. The results show that a large value of $\lambda$ reduces the FD score of both base and new classes, approaching that of zero-shot CLIP. This phenomenon indicates that the regularization in KgCoOp can prevent the model from increasing the textual divergence caused by CE loss. Correspondingly, the fine-tuned CLIP by KgCoOp preserves the same confidence level as the zero-shot CLIP. However, the fine-tuning substantially improves the accuracy of CLIP on base classes, resulting in the underconfidence issue due to the anchored confidence level. In this way, we explain why KgCoOp leads to underconfidence on base classes.
@@
Through the perspective of textual divergence, we provide a thorough explanation for the calibration tradeoff caused by different prompt-tuning methods. In addition, we present in Appendix A a theoretical justification for the relationship between textual divergence and model confidence. Ideally, we expect to maintain the excellent calibration of zero-shot CLIP after fine-tuning. In the following, we proceed by introducing our method, targeting this problem.

### v4

@@ Abstract
Confidence calibration is critical for the safe deployment of machine learning models in the real world. However, such issue in vision-language models like CLIP, particularly after fine-tuning, has not been fully addressed. In this work, we demonstrate that existing prompt tuning methods usually lead to a trade-off of calibration between base and new classes: the cross-entropy loss used in standard prompt tuning (e.g., CoOp) causes overconfidence in new classes by increasing textual label divergence, whereas regularization-based tuning (e.g., KgCoOp) maintains the confidence level but results in underconfidence in base classes due to the improved accuracy. Inspired by the observations, we introduce Dynamic Outlier Regularization (DOR) to ensure the confidence calibration on both base and new classes after fine-tuning. In particular, DOR minimizes the feature deviation of novel textual labels (instead of base classes) sampled from a large vocabulary set. In effect, DOR prevents the increase in textual divergence for new labels while easing restrictions on base classes. Extensive experiments demonstrate that DOR can notably enhance the calibration performance of current fine-tuning methods. Our code is available at https://github.com/ml-stat-Sustech/Outlier-Calibration.
Machine Learning, ICML
@@ 1 Introduction
Large pre-trained vision-language models (VLMs) like CLIP (Radford et al., 2021) have become the de facto standard in today’s zero-shot tasks including image recognition (Wortsman et al., 2022), open-vocabulary segmentation (Liang et al., 2023) and knowledge-augmented retrieval (Ming & Li, 2024). To efficiently adapt CLIP to domain-specific downstream tasks, various parameter-efficient fine-tuning (PEFT) techniques, such as prompt tuning (Zhou et al., 2022b) and adapter (Gao et al., 2024), have been proposed.  While these methods improve accuracy, they largely overlook reliability concerns such as confidence calibration in fine-tuned CLIP. Without fully understanding the miscalibration in fine-tuned CLIP, it can exacerbate safety concerns in high-stakes applications like medical diagnosis and autonomous driving.
@@
In confidence calibration, we generally expect the model’s confidence level to be consistent with its empirical accuracy (Tu et al., 2023). In the literature, zero-shot CLIP is often recognized for its excellent performance in confidence calibration (Minderer et al., 2021). Prior work (Wang et al., 2024) has shown that fine-tuning CLIP on base classes often leads to miscalibration on novel classes within the same task, where the model is expected to generalize  (Zhou et al., 2022b; Yao et al., 2023). While these studies focus on miscalibration in novel classes unseen during fine-tuning, they do not adequately address calibration issues on base classes. Overall, the community still lacks a comprehensive understanding of the fundamental cause of miscalibration in fine-tuned CLIP and effective strategies for mitigation.
@@
In this work, we first investigate how existing prompt tuning methods affect the confidence calibration of CLIP. In general, these methods can be fell into two categories: standard fine-tuning (e.g., CoOp (Zhou et al., 2022b)) and regularization-based fine-tuning (e.g., KgCoOp (Yao et al., 2023)). Our empirical analysis first reveals that both approaches struggle to maintain calibration across base and novel classes simultaneously, inevitably compromising one: CoOp exhibits overconfidence on new classes while KgCoOp provides underconfident predictions on base classes. To thoroughly understand such a phenomenon, we explain it from the perspective of textual feature divergence. In particular, CoOp increases the divergence of textual label distribution through cross-entropy loss, resulting in excessively high confidence misaligned with actual accuracy on new classes. Conversely, KgCoOp constrains the divergence increase, preserving confidence levels across both base and novel classes. However, this leads to the underconfidence issue due to the improved accuracy on base classes. This raises a key question: Can we achieve reliable calibration on both base and novel classes after fine-tuning?
@@
To tackle the above challenges, we introduce Dynamic Outlier Regularization (DOR), a regularization technique incorporated into the fine-tuning phase. The core idea of DOR is to leverage outliers to regulate the divergence of unseen textual distribution in the fine-tuned CLIP, without interfering with the vanilla fine-tuning objectives. Specifically, we first collect textual outliers from the large lexical database (e.g., WordNet (Miller, 1995)), while ensuring that the selected words do not overlap with the base classes in the fine-tuning task. During fine-tuning, we minimize the feature discrepancy of novel textual labels between the fine-tuned model and the zero-shot CLIP. To further enhance flexibility, the textual outliers are dynamically sampled from the constructed set in each epoch. By incorporating dynamic textual outliers, DOR mitigates the increase in textual divergence for novel labels while reducing constraints on base classes, improving overall calibration stability.
@@
We verify the effectiveness of DOR across 11 image classification datasets and 4 types of ImageNets with covariant shifts. Extensive experiments show that DOR can enhance the calibration of existing prompt-tuning methods (see Table 1). For instance, DOR reduces the Expected Calibration Error (ECE) by an average of 8.09% for CoOp across the 11 downstream datasets. Moreover, our method can maintain and even improve the generalization performance of those tuning methods (See Table 6 ). DOR also achieves competitive improvements in the presence of covariate shifts, further validating its robustness. Beyond prompt-based tuning, we show that this regularization criterion can be extended to visual fine-tuning methods using image outliers, highlighting its broader applicability.
@@
    We show that current prompt-tuning methods typically lead to a trade-off between base and new classes, compromising one of them. To understand such miscalibration, we provide an in-depth analysis from the perspective of textual distribution divergence.
@@
    We propose DOR, a simple yet effective regularization that ensures calibration performance on both base and new classes. Our method is compatible with existing prompt-tuning methods and can be extended to visual fine-tuning methods with image outliers.
@@
    Extensive experiments on a wide range of real-world benchmarks demonstrate that DOR is effective and easy-to-use. DOR can be easily incorporated with various prompt tuning methods.
@@ 2 Preliminaries
(a) Base
(b) New
Figure 1: Reliability diagram of fine-tuned CLIP (ViT-B/16) on StanfordCars dataset. ECE: Expected Calibration Error (lower is better). Miscalibration is depicted in pink for overconfidence and purple for underconfidence.
@@
CLIP is a visual-language model that measures the alignment between images and texts (Radford et al., 2021). Recently, CLIP has shown great potential in zero-shot inference for arbitrary classes. Let $\phi:\boldsymbol{x}\rightarrow\mathbb{R}^{d}$ and $\psi:\boldsymbol{t}\rightarrow\mathbb{R}^{d}$ denote CLIP’s image and text encoders, respectively. Given an image instance $\boldsymbol{x}$ and a text label $c$, the logit function of CLIP can be formulated as:
$L_{c}^{clip}\left(\boldsymbol{x}_{i}\right)=\tau\cdot\operatorname{sim}\left(\phi(\boldsymbol{x}),\psi(\boldsymbol{t}_{c})\right).$ (1)
Here $\boldsymbol{t}_{c}$ is derived from a hand-crafted prompt like “a photo of a {class}”, where the “{class}” is filled with the text label $c$. $\tau$ is generally set as a pre-trained constant of 100.
@@
$c^{*}=\underset{c\in\mathcal{C}}{\arg\max}\ p(c|\boldsymbol{x})=\underset{c\in\mathcal{C}}{\arg\max}\ \frac{e^{L_{c}^{clip}\left(\boldsymbol{x}\right)}}{\sum_{i=1}^{C}e^{L_{i}^{clip}\left(\boldsymbol{x}\right)}}$ (2)
where $p(c|\boldsymbol{x})$ is the predicted probability of class $c$ for the instance $\boldsymbol{x}$.
@@
To boost performance of CLIP in downstream tasks, prompt tuning methods have been proposed to efficiently fine-tune CLIP on datasets of interest (Zhou et al., 2022a, b). In particular, prompt tuning optimizes the context prompt only, without retraining the model and updating its weights. CoOp (Zhou et al., 2022b) replaces the hand-crafted textual tokens with a set of learnable textual token $\mathcal{T}=\left\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{M}\right\}$, where $M$ is the length of tokens. Thus, the output of fine-tuned CLIP is: $L_{c}^{coop}\left(\boldsymbol{x}\right)=\tau\cdot\operatorname{sim}\left(\phi(\boldsymbol{x}),\psi(\boldsymbol{t}^{\prime}_{c})\right),$ where $\boldsymbol{t}_{c}^{\prime}=\left[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3},...\boldsymbol{v}_{M},\boldsymbol{c}\right]$ and $\boldsymbol{c}$ denotes the textual embedding of class $c$. Using a few labeled samples $\mathcal{D}_{\text{ft}}=\{(\boldsymbol{x}_{i},c_{i})\}^{N}_{i=1}$, the learnable textual tokens $\mathcal{T}$ are optimized to minimize the standard cross-entropy (CE) loss $\ell_{ce}$. We refer to the classes used in fine-tuning as base classes, and the remaining labels within the same task as new or novel classes.
To enhance the generalization ability of the learnable prompt for unseen classes within the task, regularization-based methods like KgCoOp (Yao et al., 2023) introduces a regularization term to align the learned prompt to the hand-crafted prompt. The optimization of KgCoOp is:
$\begin{split}\mathcal{T}^{\star}=\underset{\mathcal{T}}{\arg\min}\Bigg{\{}&\frac{1}{N}\sum_{i=1}^{N}\ell_{ce}\left(p\left(c_{i}\mid\boldsymbol{x}_{i}\right)\right)\\ &+\lambda\cdot\frac{1}{C}\sum_{c=1}^{C}\operatorname{sim}\left(\psi\left(\boldsymbol{t}_{c}^{\prime}\right),\psi\left(\boldsymbol{t}_{c}\right)\right)\Bigg{\}}.\end{split}$ (3)
Here, the first term is the standard cross-entropy loss used in CoOp and the hyperparameter $\lambda$ is used to control the weight of regularization. With $\lambda$ = 0, KgCoOp is degraded to the original CoOp.
@@ 3.1 Empirical study on CLIP calibration
Setup
To show the miscalibration in fine-tuned CLIP, we first empirically study the calibration performance of fine-tuned VLMs. We use ViT-B-16 pre-trained by OpenAI (Radford et al., 2021) as the zero-shot classification model. In particular, we compare the zero-shot CLIP with standard prompt tuning CoOp (Zhou et al., 2022b) and regularization-based method KgCoOp (Yao et al., 2023) on StanfordCars dataset (Krause et al., 2013). We evaluate the fine-tuned CLIP under base-to-new protocol: the dataset is split into base and new classes. The model is trained only on a few examples from the base classes and evaluated on examples from both base and new classes.
Prompt tuning leads to a trade-off between base and new classes. Figure 1 illustrates the calibration performance of zero-shot CLIP, CoOp and KgCoOp on base and new classes. The results show that zero-shot CLIP achieves almost perfect calibration on all classes, while the fine-tuned models cannot maintain the calibration on base and new classes simultaneously, compromising one of them. In particular, CoOp maintains the excellent calibration on base classes but exhibits overconfidence on new classes. Instead, KgCoOp provides underconfident predictions on base classes while preserving the calibration on new classes. This motivates us to further investigate the fundamental cause of miscalibration occurring after fine-tuning.
(a) FD score
(b) Confidence
Figure 2: Results of zero-shot and fined-tuned CLIPs with different prompt tuning methods on UCF101 dataset.
@@
3.2 Understanding the miscalibration in CLIP
@@
Definition 3.1 (Feature Divergence).
@@
where $\mathcal{N}_{M}(\mathbf{z}_{i})$ denotes the set of $M$ nearest neighbors of $\mathbf{z}_{i}$ and $\operatorname{dist}\left(\cdot,\cdot\right)$ is a distance metric like cosine similarity. By averaging these similarity scores across all features, we obtain the overall FD score of a given feature set $\mathrm{FD}(\mathcal{Z})=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, which can represent the divergence of the textual distribution. To investigate the miscalibration issue, we vary the hyperparameter $\lambda$ in Eq. (3). In KgCoOp, $\lambda$ is set to 8.0, and it degenerates to CoOp when $\lambda=0$. We conduct the experiments on UCF101 (Soomro et al., 2012). We present the results in Figure 3.
@@
CoOp leads to overconfidence on new classes by increasing the textual divergence. In the analysis of Subsection 3.1 and Figure 2(b), we show that CLIP tuned by CoOp exhibits overconfidence on new classes, but keeps calibration on base classes. This is caused by the CE loss, which maximizes the posterior $p(y\mid\boldsymbol{x})$ for the ground-truth label $y$ and minimizes the probability for other labels. In other words, CE loss tends to enlarge the distance between the image feature and all textual features except the ground truth. As the image feature remains unchanged during fine-tuning, it can be translated to large distances among all textual labels, including base and new classes. This is supported by the results presented in Figure 2(a), which shows that CoOp significantly increases the FD score of textual features compared to zero-shot CLIP. Consequently, As shown in Figure 3, the gap between the maximum logit and the others widens for both base and new classes. Hence, the tuned CLIP with CE loss will make softmax predictions with high confidence on both base and new classes. It aligns with the improved accuracy on base classes, but is not consistent with the nearly unchanged accuracy on new classes. This explains why CLIP tuned by CoOp tends to be overconfident on new classes.
@@
KgCoOp anchors the confidence level by hindering the increase of textual divergence. In the previous analysis, KgCoOp leads to underconfident predictions on base classes while preserving the calibration on new classes. In Figure 2(a), we illustrate the FD scores of textual labels from the zero-shot CLIP and the tuned CLIP by KgCoOp with various $\lambda$. The results show that a large value of $\lambda$ reduces the FD score of both base and new classes, approaching that of zero-shot CLIP. This phenomenon indicates that the regularization in KgCoOp can prevent the model from increasing the textual divergence caused by CE loss. Correspondingly, the fine-tuned CLIP by KgCoOp preserves the same confidence level as the zero-shot CLIP. However, the fine-tuning substantially improves the accuracy of CLIP on base classes, resulting in the underconfidence issue due to the anchored confidence level. In this way, we explain why KgCoOp leads to underconfidence on base classes.
@@
Through the perspective of textual divergence, we provide a thorough explanation for the calibration trade-off caused by different prompt-tuning methods. We provide the comprehensive empirical results on more prompt tuning methods to thoroughly support our observationin Appendix B.1 (See Figure 5 and 6). In addition, we present a theoretical justification for the relationship between textual divergence and model confidence in Appendix C. Ideally, we expect to maintain the excellent zero-shot calibration for both base and new classes after fine-tuning. In the following, we proceed by introducing our method, targeting this problem.